# Template AASArticle with Examples: v6.3.1111Released on March, 1st, 2021

### v1

@@ 4.1 Comparison to Individual SNe Ia
Here we compare our synthetic observables from a pair of observed SNe Ia with spectra and light curves from our LTE and non-LTE calculations. To find spectral matches between our calculations and observations, an attempt was made to use SNID (Blondin & Tonry, 2007), which is intended to be used with observed spectra of SNe to classify them and estimate their redshift and epoch based on a set of templates. As the SNID algorithm removes the pseudo-continuum prior to fitting, however, we found that excellent matches determined and displayed by SNID were relatively poor matches when plotting their actual spectra due to differing spectral shapes between observation and our models. We believe that a tool that could effectively match synthetic and observed spectra while still considering for spectral shape would be beneficial to the field, especially as synthetic spectra from a variety of models improve in accuracy and the number of observed events increases.
An extensive search to determine the best matches of observables between our calculations and observed events was not undertaken due to the limited theoretical data set and scope of this work, in addition to the lack of efficient event-matching tools that the authors are aware of. Rather, we selected two samples of SNe Ia (Hicken et al. 2009 and Burrow et al. 2020), identified events therein that had similar $B$-band peak magnitudes and declines rates to our non-LTE observables, and inspected the spectra by eye to establish “matches”. We show two comparisons that involve wedges 7 and 11 from our double detonation model.
@@
Event z E(B-V) Wedge MB $\Delta$ M15,B
Obs. Model Obs. Model
SN 1999ek 0.018 0.184 11 -19.21 -19.20 1.21 1.16
SN 2007bd 0.032 0.025 7 -19.28 -19.09 1.27 0.93
Spectra of observed events are modified correcting for distance and redshift, as well as Milky Way and host galaxy reddening, using the SNooPy Python tool for SNe Ia (Burns et al., 2011) to deredden. Light curves of observed events are modified by correcting for distance, time dilation, galactic extinction and host galaxy extinction. Values for galactic extinction are drawn from Schlafly & Finkbeiner (2011) while host galaxy extinction is determined using $R_{V}=1.7$ with color excesses given by Hicken et al. (2009). We show a summary of these observations and our corresponding model wedges in Table 2. Data for our model in this table is given for only the pseudo-2D synthetic observables.
@@ 4.1.1 SN 2007bd
Figure 6: Spectrum of SN 2007bd (Blondin et al., 2012; Folatelli et al., 2013) compared with our pseudo-2D non-LTE and 2D LTE spectra for wedge 7 of our model. Spectra are labeled based on the time relative to their respective $B$-band maximum.
We show spectra of our pseudo-2D non-LTE and 2D LTE calculations along with SN 2007bd (Blondin et al., 2012; Folatelli et al., 2013) in Figure 6. Just before maximum light, the non-LTE spectrum is in modestly improved agreement with SN 2007bd, including the reproduction of an absorption feature at 4400 Å. We attribute this feature to Si iii $\lambda 4560$ which is very well-matched by the non-LTE spectrum, while the LTE spectrum shows no sign of this feature at this, or any line of sight.
At 9 days after maximum light, the non-LTE spectrum much more starkly resembles SN 2007bd. Nearly all of the observed spectrum is well-reproduced by the non-LTE spectrum at this time while the LTE shows poor overall agreement, especially blueward of $\sim 4400$. At 13.5 days post-maximum light, the agreement between the non-LTE and observed spectrum is less notable, but the LTE spectrum struggles even more than at earlier times, with reduced emission and increased line blanketing at low wavelengths associated with the lower temperature and ionization of the LTE calculation (see Figure 4).
Figure 7: Light curves of SN 2007bd (Stritzinger et al., 2011) compared with that from each of our radiative transfer calculations of wedge 7 of our model.
The light curves for SN 2007bd (Stritzinger et al., 2011) are shown in Figure 7 where they are compared to each of our radiative transfer calculations of wedge 7. The non-LTE light curves exhibit more similar shapes to SN 2007bd compared to the LTE, especially in the $I$-band where the non-LTE light curve matches the decline of the observed event 5 days after peak brightness. Additionally, while the peak brightness in $B$-band between observation and our synthetic light curves are not in complete agreement, the non-LTE light curves replicate the shallower decline from peak. All of our synthetic observables are much brighter in $U$-band at peak compared to SN 2007bd, however, the rate of decline in this band post-peak is much better captured by the non-LTE light curves.
@@ 4.1.2 SN 1999ek
Figure 8: Spectrum of SN 1999ek (Blondin et al., 2012; Silverman et al., 2012) compared with the pseudo-2D non-LTE and 2D LTE spectra of wedge 11 of our model. Spectra are labeled based on the time relative to their respective $B$-band maximum.
We show a spectral comparison between SN 1999ek (Blondin et al., 2012; Silverman et al., 2012) and both our 2D LTE and pseudo-2D non-LTE synthetic spectra for wedge 11 in Figure 8. At maximum light, the LTE and non-LTE spectra shown broadly mimic the spectrum of SN 1999ek to a similar degree. There are some differences in the two calculation’s ability to reproduce the observed spectrum however, such as in the Ca H & K line depth and shape where the non-LTE spectrum shows a local peak within the trough. On the other hand, the non-LTE spectrum shows the Si iii $\lambda 4560$ feature in its maximum light spectrum that is well-matched to SN 1999ek.
@@
Figure 9: Bolometric and UBVRI light curves for all four sets of our synthetic observables for wedge 11, compared with photometry of SN 1999ek (Krisciunas et al., 2004).
@@ 4.2 Observational Correlations
Figure 10: Phillips relation for observed SNe Ia from Hicken et al. (2009) and our four sets of calculations. Markers for our theoretical data have saturations based on their line of sight, with saturation decreasing from the northernmost line of sight. We do not show the two southernmost lines of sight for the 1D LTE and pseudo-2D non-LTE calculations on this plot due to issues with the 1D LTE calculations for these wedges as described in Section 3.1.
We compare the observed Phillips relation (Phillips, 1993; Hicken et al., 2009) to that from our single 1.0 M model across its 15 lines of sight for all four of our synthetic observable sets in Figure 10. The observational data has been adjusted for Galactic extinction and K-corrections (see Hicken et al. 2009), in addition to distance and host galaxy extinction. Both the 1D and 2D LTE calculations for this model are discrepant with the observed Phillips relation, with decline rates in the $B$-band that are appreciably larger than those seen in SNe Ia of similar peak magnitudes. Additionally, the 1D LTE calculation shows an inability to precisely recreate the relatively smooth line of sight dependence seen in the 2D LTE.
@@
In the overall sub-Chandrasekhar scenario for SNe Ia, WD mass is the main parameter for the Phillips relation, with more massive progenitors leading to brighter and slower declining events (Blondin et al., 2017b; Shen et al., 2018; Polin et al., 2019; Shen et al., 2021a). We find for our one examined model that lines of sight with brighter $B$-band peaks decline more quickly than those that have dimmer peaks when using non-LTE. In other words, the non-LTE Phillips relation for our individual model is nearly inverted from the observed Phillips relation. We emphasize that the double detonation model explored in this work is of a single mass at several lines of sight. Based on the non-trivial transformation of the observables from LTE to non-LTE (both 1D and pseudo-2D), it is challenging to precisely predict how the results from other masses, both above and below 1.0 M, will manifest on the Phillips relation. However, it may be possible for the overall theoretical Phillips relation for a suite of masses, along with the variability imposed by line of sight diversity, to reconstruct that of the observed relation. For example, the equatorial line of sight for a variety of masses may fit the general Phillips relation (resembling Figure 3 of Shen et al. 2021a), while the other lines of sight contribute to the scatter.
Figure 11: Color-magnitude plots for our four sets of synthetic observables, along with observational data from Burrow et al. (2020) and Hicken et al. (2009). Markers for our theoretical data have saturations based on their line of sight, with saturation decreasing from the northernmost line of sight. The left column of plots show observed magnitudes and colors that are corrected for Galactic extinction and distance, while the right column of plots show those that are additionally corrected for host galaxy extinction. We also include example host galaxy correction trajectories.
@@
Similar to previous works on the double detonation scenario (Polin et al., 2019; Shen et al., 2021b; Collins et al., 2022), the color determined from LTE calculations of this double detonation model are much redder than what is observed across each of the color metrics at a given peak magnitude. However, much greater agreement is seen between the non-LTE calculations and observations. Our non-LTE results are congruent with the observations that have not been attempted to be corrected for host galaxy extinction and are roughly on the boundary of the population of events that have been corrected for host galaxy extinction. Additionally, the colors of our non-LTE observables show a similar trend with peak $B$-band magnitude within the relevant magnitude range, compared to that of our LTE calculations which show a much shallower MB-(Bmax--Vmax) relation. Excluding the boundary lines of sight, we also highlight the relatively minor difference that the multidimensional effects have on the color-magnitude relation, especially for (Bmax--Vmax).
@@
Figure 12: Spectral feature comparison for our models and observed SNe Ia from Burrow et al. (2020). Both observational and theoretical values shown are calculated using Spextractor (Papadogiannakis (2019), modified by Burrow et al. (2020)). The Branch classes (Branch et al., 2006) are labeled and delineated by a dotted line in the top panel, based on approximate boundary values from Folatelli et al. (2013). In the bottom panel, the Wang classes (Wang et al., 2009) are labeled and separated by a dotted line. Spextractor did not detect a Si ii $\lambda$5972 pEW in the pseudo-2D maximum light spectra of wedge 1, so it is not included in the top panel.
@@
We show in the upper panel of Figure 12 how a few spectral indicators from our set of model observables fare in the Branch (Branch et al., 2006) classification of SNe Ia. Observed events are classified in the Branch system based on the pEW of two ubiquitous absorption features at maximum light, Si ii $\lambda$5972 and Si ii $\lambda$6355. A majority of observed SNe Ia fall within a small Si ii $\lambda$5972 pEW region, where they are sorted into shallow-silicon (SS), core-normal (CN), and broad-line (BL) classes based on Si ii $\lambda$6355 pEW in an increasing order. We note that these classes are relatively continuous and are not necessarily considered to be distinct groups of events. The 1D and 2D LTE calculations of our one model straddle the boundary between these groups and the cool (CL) class of SNe Ia, specifically at a region in this parameter space that has a dearth of observed events. In the non-LTE treatments of the synthetic observables, however, the Si ii $\lambda$5972 pEWs are consistently lower at all lines of sight and are brought into much better agreement with a portion of observed SNe Ia for these spectral indicators. The non-LTE spectra span three Branch classes, with the southern, equatorial, and northern lines of sight representing the SS, CN, and BL classes, respectively. This may be an indication that these identified classes of SNe Ia can be at least partially attributed to line of sight. We also find that the extent of the Si ii $\lambda$6355 pEW is wider in our non-LTE treatment of the synthetic observables compared to LTE, nearly spanning the breadth from that of observed SNe Ia.
@@
In the lower panel of Figure 12, we show the pEW and velocity of Si ii $\lambda$6355 for our models and observed SNe Ia. The calculated Si ii $\lambda$6355 velocity is fairly consistent between the non-LTE and LTE treatments for each line of sight, with more northern lines of sight showing higher velocities. We delineate this plot based on the Wang classification (Wang et al., 2009) that categorizes observed events with Si ii $\lambda$6355 velocity above $\sim 11,000$ km s-1 as high velocity (HV). Wang et al. (2009) shows that HV events are redder on average, either intrinsically or otherwise. Interestingly, our model spectra that could be classified as HV occur at northern lines of sight, which also see redder colors (see Figure 11), another indication that some spectral classes may be able to be explained with line of sight effects. The normal and HV Wang classes on average do not have distinctly different peak magnitudes as our individual model does across lines of sight, however. The characteristics shown in Figure 12 are modestly more impacted by dimensionality than what was shown previously for the Phillips relation and color-magnitude, likely due to the increased sensitivity of these parameters.

### v2

@@ 4.1 Comparison to Individual SNe Ia
Here we compare our synthetic observables from a pair of observed SNe Ia with spectra and light curves from our LTE and non-LTE calculations. To find spectral matches between our calculations and observations, an attempt was made to use SNID (S. Blondin & J. L. Tonry, 2007), which is intended to be used with observed spectra of SNe to classify them and estimate their redshift and epoch based on a set of templates. As the SNID algorithm removes the pseudo-continuum prior to fitting, however, we found that excellent matches determined and displayed by SNID were relatively poor matches when plotting their actual spectra due to differing spectral shapes between observation and our models. We believe that a tool that could effectively match synthetic and observed spectra while still considering for spectral shape would be beneficial to the field, especially as synthetic spectra from a variety of models improve in accuracy and the number of observed events increases.
An extensive search to determine the best matches of observables between our calculations and observed events was not undertaken due to the limited theoretical data set and scope of this work, in addition to the lack of efficient event-matching tools that the authors are aware of. Rather, we selected two samples of SNe Ia (M. Hicken et al. 2009 and A. Burrow et al. 2020), identified events therein that had similar $B$-band peak magnitudes and declines rates to our non-LTE observables, and inspected the spectra by eye to establish “matches”. We additionally compare to SN 2011fe. In the following subsections, we compare the observables from these normal SNe Ia with our wedges that are most similar.
@@
Event z E(B-V)host Wedge MB,max $\Delta$M15,B
mag mag mag
Observed ModelaaPseudo-2D non-LTE Observed ModelaaPseudo-2D non-LTE
SN 2011fe 0.001 12 $-19.21\pm 0.15$bbM. W. Richmond & H. A. Smith 2012 $-19.23$ $1.21\pm 0.03$bbM. W. Richmond & H. A. Smith 2012 1.37
SN 1999ek 0.018 0.184 11 $-19.21\pm 0.47$ccM. Hicken et al. 2009 $-19.22$ $1.21\pm 0.37$ccM. Hicken et al. 2009 1.22
SN 2007bd 0.032 0.025 7 $-19.28\pm 0.10$ddA. Burrow et al. 2020 $-19.10$ $1.24\pm 0.01$eeM. D. Stritzinger et al. 2011 1.02
Spectra of observed events are modified correcting for distance and redshift, as well as Milky Way and host galaxy reddening, using the SNooPy Python tool for SNe Ia (C. R. Burns et al., 2011) to deredden. Light curves of observed events are modified by correcting for distance, time dilation, galactic extinction and host galaxy extinction (we do not consider host galaxy extinction for SN 2011fe). Values for galactic extinction are drawn from E. F. Schlafly & D. P. Finkbeiner (2011) while host galaxy extinction is determined using $R_{V}=1.7$ with color excesses given by M. Hicken et al. (2009). We show a summary of these observations and our corresponding model wedges in Table 2. Data for our model in this table is given for only the pseudo-2D synthetic observables.
4.1.1 SN 2011fe
Figure 6: Spectrum of SN 2011fe (R. Pereira et al., 2013; P. A. Mazzali et al., 2014) compared with our 1D non-LTE and 2D LTE spectra for wedge 12 of our model. The cyan shading indicates the difference between 1D non-LTE and pseudo-2D non-LTE spectra at the given time. Spectra are labeled based on the time relative to their respective $B$-band maximum.
Spectra of SN 2011fe (R. Pereira et al., 2013; P. A. Mazzali et al., 2014) are compared with our 2D LTE and 1D non-LTE results for wedge 12 of our model in Figure 6. We additionally show our pseudo-2D non-LTE spectra (cyan) as a deviation from the the 1D non-LTE spectra, which gives an indication of where multidimensional effects may apply. We emphasize that this deviation shown in Figure 6, and the following spectral comparisons, display the differences between the 1D non-LTE and pseudo-2D non-LTE results with consideration for a change of the time of $B$-band maximum between sets of synthetic observables (see Figure 2 for an illustration of the multidimensional spectral differences at the same post-explosion times).
Before and at maximum light, SN 2011fe is well-reproduced by the 1D non-LTE spectra at wavelengths blueward of the Ca ii H&K feature, including the Si iii $\lambda 4560$ absorption feature. In this same region, the 2D LTE spectra fails to reproduce some features of SN 2011fe, in addition to having modestly higher emission in the optical. Much of the pseudo-2D non-LTE spectra deviates from the 1D non-LTE here, however, suggesting that the strong agreement between SN 2011fe and the 1D non-LTE spectra may be driven by the reduced multidimensionality of the calculation. Additionally, the UV spectra of SN 2011fe is not well-replicated by any of our synthetic observables at maximum light.
By 7.5 days, the disparity between the LTE and non-LTE agreement with SN 2011fe is even greater, in addition to the multidimensional correction being lower. At this and later times, non-LTE spectra more accurately reproduce the near-UV features of SN 2011fe, in addition to the waning Ca ii H&K feature. The non-LTE spectra shows improved agreement in the optical, with a lack of suppression at $\sim 4250$Å and a S ii $\lambda 5640$ absorption feature that persists longer beyond maximum light. The very near IR spectra is also much improved in the non-LTE calculations, including the Ca ii triplet at $\sim 8500$Å.
Figure 7: Light curves of SN 2011fe (U. Munari et al., 2013; D. Y. Tsvetkov et al., 2013) compared with that from each of our radiative transfer calculations of wedge 12 of our model.
We compare the photometry of SN 2011fe (U. Munari et al., 2013; D. Y. Tsvetkov et al., 2013) with each set of our synthetic observables for wedge 12 in Figure 7. While the shape of the pseudo-bolometric light curve of SN 2011fe is not well fit by our model constraint, the non-LTE $UBVRI$ light curves show a strong agreement with observation. Most notably, the 1D non-LTE $U$- and $B$-band are remarkably consistent with that from SN 2011fe, which is a significant departure from the LTE result. There is a modest multidimensional correction for the $B$-band at max light for this wedge, however. The $V$- and $R$-bands also show moderately improved agreement with SN 2011fe under non-LTE.
@@ 4.1.2 SN 2007bd
Figure 8: Spectrum of SN 2007bd (S. Blondin et al., 2012; G. Folatelli et al., 2013) compared with our pseudo-2D non-LTE and 2D LTE spectra for wedge 7 of our model. Spectra are labeled based on the time relative to their respective $B$-band maximum.
We show spectra of our pseudo-2D non-LTE and 2D LTE calculations along with SN 2007bd (S. Blondin et al., 2012; G. Folatelli et al., 2013) in Figure 8. Just before maximum light, the non-LTE spectrum is in modestly improved agreement with SN 2007bd, including the reproduction of Si iii $\lambda 4560$ at $\sim 4400$. At 9 days after maximum light, the non-LTE spectra continues to better resemble SN 2007bd. Most of the observed spectrum is well-reproduced by the non-LTE spectrum at this time while the LTE shows poor overall agreement, especially blueward of $\sim 4400$. At 13.5 days post-maximum light, the agreement between the non-LTE and observed spectrum is less notable, but the LTE spectrum struggles even more than at earlier times, with reduced emission and increased line blanketing at low wavelengths associated with the lower temperature and ionization of the LTE calculation (see Figure 4).
Figure 9: Light curves of SN 2007bd (M. D. Stritzinger et al., 2011) compared with that from each of our radiative transfer calculations of wedge 7 of our model.
The light curves for SN 2007bd (M. D. Stritzinger et al., 2011) are shown in Figure 9 where they are compared to each of our radiative transfer calculations of wedge 7. The non-LTE light curves exhibit more similar shapes to SN 2007bd compared to the LTE, especially in the $I$-band where the non-LTE light curve matches the decline of the observed event 5 days after peak brightness. Additionally, while the peak brightness in $B$-band between observation and our synthetic light curves are not in complete agreement, the non-LTE light curves replicate the shallower decline from peak. All of our synthetic observables are much brighter in $U$-band at peak compared to SN 2007bd, however, the rate of decline in this band post-peak is much better captured by the non-LTE light curves.
@@ 4.1.3 SN 1999ek
Figure 10: Spectrum of SN 1999ek (S. Blondin et al., 2012; J. M. Silverman et al., 2012) compared with the pseudo-2D non-LTE and 2D LTE spectra of wedge 11 of our model. Spectra are labeled based on the time relative to their respective $B$-band maximum.
We show a spectral comparison between SN 1999ek (S. Blondin et al., 2012; J. M. Silverman et al., 2012) and both our 2D LTE and pseudo-2D non-LTE synthetic spectra for wedge 11 in Figure 10. At maximum light, the LTE and non-LTE spectra shown broadly mimic the spectrum of SN 1999ek to a similar degree. There are some differences in the two calculation’s ability to reproduce the observed spectrum however, such as in the Ca ii H&K line depth and shape where the non-LTE spectrum shows a local peak within the trough. On the other hand, the non-LTE spectrum shows the Si iii $\lambda 4560$ feature in its maximum light spectrum that is well-matched to SN 1999ek.
@@
Figure 11: Bolometric and UBVRI light curves for all four sets of our synthetic observables for wedge 11, compared with photometry of SN 1999ek (K. Krisciunas et al., 2004).
@@ 4.2 Observational Correlations
Figure 12: Phillips relation for observed SNe Ia from M. Hicken et al. (2009) and our four sets of calculations. Markers for our theoretical data have saturations based on their line of sight, with saturation decreasing from the northernmost line of sight.
We compare the observed Phillips relation (M. M. Phillips, 1993; M. Hicken et al., 2009) to that from our single 1.0 M model across its 15 lines of sight for all four of our synthetic observable sets in Figure 12. The observational data has been adjusted for Galactic extinction and K-corrections (see M. Hicken et al. 2009), in addition to distance and host galaxy extinction. Both the 1D and 2D LTE calculations for this model are discrepant with the observed Phillips relation, with decline rates in the $B$-band that are appreciably larger than those seen in SNe Ia of similar peak magnitudes. Additionally, the 1D LTE calculation shows an inability to precisely recreate the relatively smooth line of sight dependence seen in the 2D LTE.
@@
In the overall sub-Chandrasekhar scenario for SNe Ia, WD mass is the main parameter for the Phillips relation, with more massive progenitors leading to brighter and slower declining events (S. Blondin et al., 2017; K. J. Shen et al., 2018; A. Polin et al., 2019; K. J. Shen et al., 2021a). We find for our one examined model that lines of sight with brighter $B$-band peaks decline more quickly than those that have dimmer peaks when using non-LTE. In other words, the non-LTE Phillips relation for our individual model is nearly inverted from the observed Phillips relation. We emphasize that the double detonation model explored in this work is of a single mass at several lines of sight. Based on the non-trivial transformation of the observables from LTE to non-LTE (both 1D and pseudo-2D), it is challenging to precisely predict how the results from other masses, both above and below 1.0 M, will manifest on the Phillips relation. However, it may be possible for the overall theoretical Phillips relation for a suite of masses, along with the variability imposed by line of sight diversity, to reconstruct that of the observed relation. For example, the equatorial line of sight for a variety of masses may fit the general Phillips relation (resembling Figure 3 of K. J. Shen et al. 2021a), while the other lines of sight contribute to the scatter. This line of sight-induced scatter effect has previously been demonstrated for the double detonation (K. J. Shen et al., 2021b; C. E. Collins et al., 2022; J. M. Pollin et al., 2024), as well as the delayed detonation scenario for SNe Ia (S. A. Sim et al., 2013).
Figure 13: Color-magnitude plots for our four sets of synthetic observables, along with observational data from A. Burrow et al. (2020) and M. Hicken et al. (2009). Markers for our theoretical data have saturations based on their line of sight, with saturation decreasing from the northernmost line of sight. The left column of plots show observed magnitudes and colors that are corrected for Galactic extinction and distance, while the right column of plots show those that are additionally corrected for host galaxy extinction. We also include example host galaxy correction trajectories.
@@
Similar to previous works on the double detonation scenario (A. Polin et al., 2019; K. J. Shen et al., 2021b; C. E. Collins et al., 2022), the color determined from LTE calculations of this double detonation model are much redder than what is observed across each of the color metrics at a given peak magnitude. However, much greater agreement is seen between the non-LTE calculations and observations. Our non-LTE results are congruent with the observations that have not been attempted to be corrected for host galaxy extinction and are roughly on the boundary of the population of events that have been corrected for host galaxy extinction. Additionally, the colors of our non-LTE observables show a similar trend with peak $B$-band magnitude within the relevant magnitude range, compared to that of our LTE calculations which show a much shallower MB-(Bmax--Vmax) relation. Excluding the boundary lines of sight, we also highlight the relatively minor difference that the multidimensional effects have on the color-magnitude relation, especially for (Bmax--Vmax).
@@
Figure 14: Spectral feature comparison for our models and observed SNe Ia from A. Burrow et al. (2020). Both observational and theoretical values shown are calculated using Spextractor (S. Papadogiannakis (2019), modified by A. Burrow et al. (2020)). The Branch classes (D. Branch et al., 2006) are labeled and delineated by a dotted line in the top panel, based on approximate boundary values from G. Folatelli et al. (2013). In the bottom panel, the Wang classes (X. Wang et al., 2009) are labeled and separated by a dotted line. Spextractor did not detect a Si ii $\lambda$5972 pEW in the pseudo-2D maximum light spectra of wedge 1, so it is integrated manually.
@@
We show in the upper panel of Figure 14 how a few spectral indicators from our set of model observables fare in the Branch (D. Branch et al., 2006) classification of SNe Ia. Observed events are classified in the Branch system based on the pEW of two ubiquitous absorption features at maximum light, Si ii $\lambda$5972 and Si ii $\lambda$6355. A majority of observed SNe Ia fall within a small Si ii $\lambda$5972 pEW region, where they are sorted into shallow-silicon (SS), core-normal (CN), and broad-line (BL) classes based on Si ii $\lambda$6355 pEW in an increasing order. We note that these classes are relatively continuous and are not necessarily considered to be distinct groups of events. The 1D and 2D LTE calculations of our one model straddle the boundary between these groups and the cool (CL) class of SNe Ia, specifically at a region in this parameter space that has a dearth of observed events. In the non-LTE treatments of the synthetic observables, however, the Si ii $\lambda$5972 pEWs are consistently lower at all lines of sight and are brought into much better agreement with a portion of observed SNe Ia for these spectral indicators. The non-LTE spectra span three Branch classes, with the southern, equatorial, and northern lines of sight representing the SS, CN, and BL classes, respectively. This may be an indication that these identified classes of SNe Ia can be at least partially attributed to line of sight. We also find that the extent of the Si ii $\lambda$6355 pEW is wider in our non-LTE treatment of the synthetic observables compared to LTE, nearly spanning the breadth from that of observed SNe Ia.
@@
In the lower panel of Figure 14, we show the pEW and velocity of Si ii $\lambda$6355 for our models and observed SNe Ia. The calculated Si ii $\lambda$6355 velocity is fairly consistent between the non-LTE and LTE treatments for each line of sight, with more northern lines of sight showing higher velocities. We delineate this plot based on the Wang classification (X. Wang et al., 2009) that categorizes observed events with Si ii $\lambda$6355 velocity above $\sim 11,000$ km s-1 as high velocity (HV). X. Wang et al. (2009) shows that HV events are redder on average, either intrinsically or otherwise. Interestingly, our model spectra that could be classified as HV occur at northern lines of sight, which also see redder colors (see Figure 13), another indication that some spectral classes may be able to be explained with line of sight effects. The normal and HV Wang classes on average do not have distinctly different peak magnitudes as our individual model does across lines of sight, however. The characteristics shown in Figure 14 are modestly more impacted by dimensionality than what was shown previously for the Phillips relation and color-magnitude, likely due to the increased sensitivity of these parameters.